# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ Appendix B Additional Results
Lastly, Tables 8 and 9 contain results for our XNLI setup where the base model is fine-tuned from the pretrained mBERT with Standard FL using full fine-tuning and is used to complement results and findings in Section 4.2. Similarly, Tables 10 and 11 complement the results and findings of our Fed-Aya setup described in Section 4.3.
Figure 5: $\bm{\lambda}$ distance among languages in our MasakhaNEWS setup. Each block shows the log-scale normalized average Euclidean distances between all pairs of clients’ $\bm{\lambda}$ for two languages (see text). The smaller the distance, the more similar $\bm{\lambda}$ is.
@@
$\mathbf{r}$ Approach bg hi es el vi tr de ur en zh th sw ar fr ru Wins
2 LoRA 47.47±0.19 45.93±0.34 50.80±0.16 50.80±0.16 50.80±0.28 48.80±0.59 50.07±0.66 49.67±0.47 53.13±0.41 50.00±0.28 45.47±0.41 44.33±0.09 45.33±0.25 51.00±0.16 48.33±0.74 0
AdaLoRA 45.73±0.09 44.40±0.00 49.00±0.28 49.13±0.19 48.00±0.16 47.07±0.34 48.27±0.09 47.87±0.09 50.93±0.09 48.20±0.16 43.80±0.16 43.00±0.00 44.20±0.16 48.87±0.09 46.27±0.25 0
BayesTune-LoRA 45.67±0.19 44.40±0.00 48.33±0.25 48.80±0.16 48.13±0.19 46.80±0.28 48.13±0.09 47.80±0.00 50.87±0.19 48.00±0.00 43.53±0.25 42.13±0.19 44.13±0.09 48.93±0.09 45.67±0.09 0
FedL2P 66.67±0.90 69.47±0.77 74.33±0.93 70.73±1.00 71.27±0.82 71.27±0.57 72.27±1.32 73.20±0.28 75.27±0.81 68.27±1.32 67.93±0.75 73.47±1.31 71.47±0.94 72.80±0.28 73.07±0.25 0
FedP2EFT 71.73±0.41 72.33±0.25 75.40±0.59 73.73±0.84 74.80±0.43 74.73±0.41 75.00±0.71 74.33±0.57 75.47±0.19 70.53±0.52 71.33±0.09 73.73±0.41 75.60±0.57 74.60±0.59 74.13±0.68 15
4 LoRA 49.40±0.16 50.00±0.86 54.47±0.62 53.33±0.41 53.00±0.28 51.53±0.66 53.27±0.62 52.93±1.09 56.27±0.47 52.20±0.43 49.07±0.62 48.80±0.00 49.20±0.33 54.27±0.34 51.73±1.11 0
AdaLoRA 45.87±0.19 44.33±0.09 48.60±0.16 48.93±0.09 48.27±0.34 46.87±0.09 48.47±0.09 47.80±0.00 51.27±0.09 48.07±0.09 43.67±0.09 42.73±0.25 44.20±0.28 48.73±0.09 45.93±0.19 0
BayesTune-LoRA 45.60±0.00 44.33±0.09 48.87±0.19 49.13±0.09 48.00±0.28 47.00±0.33 48.47±0.09 47.80±0.16 51.13±0.09 48.13±0.09 43.73±0.19 42.40±0.16 44.20±0.00 48.93±0.09 46.00±0.16 0
FedL2P 67.40±1.42 69.73±1.37 74.47±0.34 70.20±0.85 71.53±0.94 71.27±0.90 72.73±0.94 73.07±0.34 75.27±0.47 68.73±1.16 68.27±0.47 73.73±0.66 71.47±1.64 73.47±0.50 72.87±0.52 0
FedP2EFT 72.80±0.33 72.13±0.41 75.40±0.28 74.27±0.34 74.93±0.09 75.20±0.28 75.80±0.16 75.07±0.25 75.93±0.25 71.60±0.16 70.80±0.28 73.07±0.25 75.53±0.62 75.87±0.34 74.87±0.19 15
8 LoRA 55.33±1.60 54.13±0.09 59.93±0.50 58.20±1.28 58.07±0.25 56.53±0.25 59.73±0.09 58.47±1.33 63.40±0.71 57.13±0.34 56.13±0.19 56.13±0.50 55.93±0.09 59.20±0.33 58.20±2.01 0
AdaLoRA 45.80±0.00 44.40±0.00 48.47±0.09 49.00±0.16 48.13±0.09 46.93±0.09 48.27±0.09 48.00±0.16 50.80±0.16 48.07±0.09 43.73±0.09 42.40±0.16 44.20±0.16 48.93±0.09 45.93±0.09 0
BayesTune-LoRA 45.87±0.25 44.33±0.09 49.13±0.47 49.27±0.09 48.20±0.00 47.07±0.09 48.60±0.16 47.80±0.00 51.00±0.28 48.60±0.16 44.13±0.09 42.93±0.09 44.40±0.00 49.20±0.43 46.20±0.16 0
FedL2P 64.73±1.27 66.07±2.19 72.13±1.80 67.60±1.42 68.80±1.85 68.47±1.89 69.40±1.99 70.93±2.22 73.47±1.18 65.93±1.86 65.73±1.98 70.73±2.96 67.73±1.93 70.53±1.93 70.47±1.89 0
FedP2EFT 73.93±0.34 70.67±0.09 75.87±0.19 73.80±0.16 74.33±0.47 75.60±0.16 74.93±0.09 74.27±0.19 76.47±0.09 72.53±0.52 70.67±0.09 71.67±0.34 76.87±0.25 75.87±0.25 75.00±0.57 15
16 LoRA 63.93±1.46 64.20±0.00 70.40±0.16 67.07±1.32 68.53±0.25 66.07±0.66 68.13±0.25 69.67±0.62 72.47±0.90 64.67±1.24 65.47±1.11 69.87±0.90 67.20±0.16 68.73±0.09 68.00±0.85 0
AdaLoRA 45.80±0.16 44.40±0.00 48.73±0.25 48.93±0.19 48.07±0.19 47.00±0.33 48.27±0.09 47.73±0.09 50.93±0.19 48.20±0.16 43.47±0.09 42.53±0.09 44.20±0.00 48.93±0.09 45.80±0.00 0
BayesTune-LoRA 46.27±0.09 44.93±0.19 49.67±0.25 49.60±0.00 48.60±0.16 47.47±0.19 49.00±0.00 48.00±0.00 51.47±0.19 49.00±0.16 44.67±0.09 43.07±0.25 44.33±0.25 49.13±0.52 46.67±0.34 0
FedL2P 61.53±2.88 62.33±3.44 68.73±3.60 65.07±1.88 65.13±3.10 65.07±3.20 65.80±3.59 67.73±2.87 71.33±2.13 62.73±2.78 62.27±3.27 67.20±3.92 63.93±3.77 68.07±2.88 67.20±3.43 0
FedP2EFT 73.87±0.09 70.80±0.16 75.87±0.09 73.73±0.68 74.93±0.25 75.33±0.25 74.40±0.16 74.13±0.19 76.53±0.19 72.40±0.28 70.80±0.28 71.13±0.62 76.20±0.43 75.87±0.09 75.13±0.09 15

### v2

@@ Appendix E Additional Results
$\mathbf{r}$ 2 4
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 54.1±0.1 53.3±0.1 53.4±0.0 64.7±1.1 70.7±0.8 56.7±0.3 53.3±0.1 53.3±0.1 66.5±1.1 71.3±0.3
hi 55.8±0.0 54.5±0.1 54.4±0.0 64.2±2.8 70.1±1.0 59.1±0.2 54.4±0.2 54.5±0.2 65.4±1.8 70.1±1.1
es 57.1±0.1 55.3±0.1 55.4±0.0 67.9±1.3 73.9±1.5 59.1±0.4 55.3±0.1 55.4±0.2 69.9±1.9 73.3±2.3
el 55.1±0.3 52.8±0.3 53.1±0.1 68.5±1.1 73.5±0.2 57.3±0.2 53.2±0.0 53.0±0.2 70.5±0.9 73.3±0.7
vi 56.3±0.1 54.1±0.1 54.0±0.2 68.6±0.2 73.4±0.6 59.0±0.3 54.1±0.2 54.5±0.2 70.1±0.3 72.6±0.3
tr 54.3±0.1 52.9±0.1 52.9±0.1 65.2±4.0 71.9±2.1 56.9±0.2 52.9±0.1 52.9±0.1 66.9±4.1 71.9±2.3
de 55.1±0.1 53.9±0.1 54.1±0.1 67.3±0.7 73.3±0.6 57.4±0.6 54.0±0.0 54.0±0.0 69.7±0.9 74.6±0.2
ur 53.6±0.2 52.9±0.2 53.2±0.0 67.0±0.4 71.9±0.6 55.9±0.1 52.9±0.1 53.1±0.1 68.2±0.6 72.9±0.4
en 61.5±0.1 60.4±0.2 60.0±0.0 71.5±0.3 74.5±0.5 63.5±0.2 60.3±0.2 60.6±0.2 72.8±0.2 75.0±0.3
zh 54.8±0.0 53.1±0.1 53.2±0.0 64.9±0.5 73.8±0.4 57.3±0.1 53.1±0.1 53.1±0.1 67.4±1.2 74.3±1.2
th 50.7±0.1 49.4±0.2 49.3±0.1 61.7±1.7 65.6±3.0 52.5±0.3 49.2±0.2 49.3±0.2 62.6±2.0 66.1±3.2
sw 52.5±0.1 50.7±0.1 50.5±0.1 64.6±2.8 69.9±1.2 54.5±0.2 50.6±0.2 51.1±0.1 65.7±2.9 68.1±0.3
ar 53.4±0.2 51.4±0.2 51.5±0.1 65.8±2.0 74.4±0.5 56.4±0.3 51.3±0.1 51.5±0.1 67.7±1.3 75.1±1.6
fr 53.1±0.1 52.8±0.2 52.6±0.2 63.4±0.0 71.3±1.5 55.9±0.1 52.6±0.2 52.6±0.0 66.5±1.1 72.2±0.3
ru 55.3±0.1 54.2±0.2 54.0±0.0 67.4±0.4 72.7±1.9 58.4±0.3 54.0±0.0 54.5±0.2 69.0±0.4 73.5±1.6
$\mathbf{r}$ 8 16
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 60.6±0.3 53.3±0.1 53.3±0.1 66.4±0.6 70.5±0.2 67.1±0.3 53.4±0.0 53.5±0.1 68.0±1.0 69.8±0.2
hi 62.5±0.1 54.5±0.2 54.4±0.0 65.5±1.1 69.3±0.8 67.8±0.0 54.4±0.0 54.7±0.1 65.9±1.3 69.3±1.0
es 63.7±0.1 55.3±0.1 55.5±0.2 70.5±3.5 73.3±1.8 69.5±0.1 55.4±0.0 55.8±0.0 70.2±3.1 73.5±1.5
el 62.6±0.3 52.8±0.3 53.2±0.0 70.1±0.5 71.3±1.4 71.3±0.2 52.6±0.2 53.4±0.0 71.7±0.8 70.4±1.8
vi 64.7±0.1 54.3±0.2 54.7±0.1 70.5±1.7 71.3±0.5 69.2±0.0 54.0±0.0 54.7±0.1 71.2±1.4 71.3±0.6
tr 61.9±0.4 53.0±0.0 53.0±0.0 68.2±2.2 70.5±2.1 68.1±0.4 52.9±0.1 53.1±0.1 69.6±1.7 70.7±2.9
de 62.5±0.1 53.9±0.1 54.1±0.1 69.6±0.4 75.3±0.4 69.0±0.3 53.9±0.1 54.5±0.2 71.2±1.0 74.8±0.2
ur 60.3±0.2 52.9±0.1 53.1±0.2 67.5±1.3 73.1±0.8 68.7±0.2 52.9±0.1 53.3±0.1 69.0±1.0 73.2±1.1
en 67.3±0.2 60.1±0.3 60.6±0.0 72.6±0.6 74.4±0.8 71.5±0.2 60.3±0.4 60.7±0.1 73.8±0.8 74.5±1.2
zh 61.4±0.2 53.0±0.0 53.5±0.1 67.1±1.3 73.9±1.1 68.0±0.2 53.0±0.0 53.4±0.2 69.1±1.5 73.5±0.8
th 56.5±0.1 49.2±0.2 49.5±0.1 61.8±1.0 66.7±3.8 62.8±0.0 49.3±0.1 49.4±0.0 64.0±1.1 65.9±2.6
sw 59.3±0.1 50.3±0.2 51.3±0.3 65.6±1.8 68.1±0.9 67.3±0.2 50.3±0.1 51.7±0.1 67.8±1.8 68.6±1.7
ar 61.3±0.2 51.2±0.2 51.7±0.2 68.2±1.0 74.7±1.2 66.8±0.2 51.2±0.3 52.1±0.2 69.3±1.2 73.7±1.8
fr 58.9±0.1 52.7±0.4 52.9±0.1 67.2±1.4 71.9±0.2 65.1±0.2 52.8±0.3 53.1±0.1 68.5±1.3 70.7±0.2
ru 63.1±0.2 53.9±0.1 54.4±0.2 69.1±1.7 72.9±1.0 67.3±0.1 53.8±0.2 54.8±0.0 69.9±1.9 72.8±0.7
Additional Quantitative Results. We show extended results for 1) our MasakhaNEWS setup in Tables 7 and 8, 2) our XNLI setup with FedDPA-T used to pretrain the base model in Table 9, and 3) our Fed-Aya setup with off-the-shelf Llama3 in Tables 10 and 11 for seen and unseen clients respectively.
We also present additional results to supplement the results in the main paper, namely 1) our XNLI setup with Standard FL (full fine-tuning) used to pretrain the base model in Tables 13 and 14, 2) our XNLI setup to show that FedP2EFT is complementary with other pFL methods in Table 12, and 3) our Fed-Aya setup with MobileLLaMA as the base model in Tables 15 and 16.
$\mathbf{r}$ 2 4
Approach LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 47.5±0.2 45.7±0.1 45.7±0.2 66.7±0.9 71.7±0.4 49.4±0.2 45.9±0.2 45.6±0.0 67.4±1.4 72.8±0.3
hi 45.9±0.3 44.4±0.0 44.4±0.0 69.5±0.8 72.3±0.2 50.0±0.9 44.3±0.1 44.3±0.1 69.7±1.4 72.1±0.4
es 50.8±0.2 49.0±0.3 48.3±0.2 74.3±0.9 75.4±0.6 54.5±0.6 48.6±0.2 48.9±0.2 74.5±0.3 75.4±0.3
el 50.8±0.2 49.1±0.2 48.8±0.2 70.7±1.0 73.7±0.8 53.3±0.4 48.9±0.1 49.1±0.1 70.2±0.8 74.3±0.3
vi 50.8±0.3 48.0±0.2 48.1±0.2 71.3±0.8 74.8±0.4 53.0±0.3 48.3±0.3 48.0±0.3 71.5±0.9 74.9±0.1
tr 48.8±0.6 47.1±0.3 46.8±0.3 71.3±0.6 74.7±0.4 51.5±0.7 46.9±0.1 47.0±0.3 71.3±0.9 75.2±0.3
de 50.1±0.7 48.3±0.1 48.1±0.1 72.3±1.3 75.0±0.7 53.3±0.6 48.5±0.1 48.5±0.1 72.7±0.9 75.8±0.2
ur 49.7±0.5 47.9±0.1 47.8±0.0 73.2±0.3 74.3±0.6 52.9±1.1 47.8±0.0 47.8±0.2 73.1±0.3 75.1±0.2
en 53.1±0.4 50.9±0.1 50.9±0.2 75.3±0.8 75.5±0.2 56.3±0.5 51.3±0.1 51.1±0.1 75.3±0.5 75.9±0.2
zh 50.0±0.3 48.2±0.2 48.0±0.0 68.3±1.3 70.5±0.5 52.2±0.4 48.1±0.1 48.1±0.1 68.7±1.2 71.6±0.2
th 45.5±0.4 43.8±0.2 43.5±0.2 67.9±0.8 71.3±0.1 49.1±0.6 43.7±0.1 43.7±0.2 68.3±0.5 70.8±0.3
sw 44.3±0.1 43.0±0.0 42.1±0.2 73.5±1.3 73.7±0.4 48.8±0.0 42.7±0.2 42.4±0.2 73.7±0.7 73.1±0.2
ar 45.3±0.2 44.2±0.2 44.1±0.1 71.5±0.9 75.6±0.6 49.2±0.3 44.2±0.3 44.2±0.0 71.5±1.6 75.5±0.6
fr 51.0±0.2 48.9±0.1 48.9±0.1 72.8±0.3 74.6±0.6 54.3±0.3 48.7±0.1 48.9±0.1 73.5±0.5 75.9±0.3
ru 48.3±0.7 46.3±0.2 45.7±0.1 73.1±0.2 74.1±0.7 51.7±1.1 45.9±0.2 46.0±0.2 72.9±0.5 74.9±0.2
$\mathbf{r}$ 8 16
Approach LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 55.3±1.6 45.8±0.0 45.9±0.2 64.7±1.3 73.9±0.3 63.9±1.5 45.8±0.2 46.3±0.1 61.5±2.9 73.9±0.1
hi 54.1±0.1 44.4±0.0 44.3±0.1 66.1±2.2 70.7±0.1 64.2±0.0 44.4±0.0 44.9±0.2 62.3±3.4 70.8±0.2
es 59.9±0.5 48.5±0.1 49.1±0.5 72.1±1.8 75.9±0.2 70.4±0.2 48.7±0.2 49.7±0.2 68.7±3.6 75.9±0.1
el 58.2±1.3 49.0±0.2 49.3±0.1 67.6±1.4 73.8±0.2 67.1±1.3 48.9±0.2 49.6±0.0 65.1±1.9 73.7±0.7
vi 58.1±0.2 48.1±0.1 48.2±0.0 68.8±1.9 74.3±0.5 68.5±0.2 48.1±0.2 48.6±0.2 65.1±3.1 74.9±0.2
tr 56.5±0.2 46.9±0.1 47.1±0.1 68.5±1.9 75.6±0.2 66.1±0.7 47.0±0.3 47.5±0.2 65.1±3.2 75.3±0.2
de 59.7±0.1 48.3±0.1 48.6±0.2 69.4±2.0 74.9±0.1 68.1±0.2 48.3±0.1 49.0±0.0 65.8±3.6 74.4±0.2
ur 58.5±1.3 48.0±0.2 47.8±0.0 70.9±2.2 74.3±0.2 69.7±0.6 47.7±0.1 48.0±0.0 67.7±2.9 74.1±0.2
en 63.4±0.7 50.8±0.2 51.0±0.3 73.5±1.2 76.5±0.1 72.5±0.9 50.9±0.2 51.5±0.2 71.3±2.1 76.5±0.2
zh 57.1±0.3 48.1±0.1 48.6±0.2 65.9±1.9 72.5±0.5 64.7±1.2 48.2±0.2 49.0±0.2 62.7±2.8 72.4±0.3
th 56.1±0.2 43.7±0.1 44.1±0.1 65.7±2.0 70.7±0.1 65.5±1.1 43.5±0.1 44.7±0.1 62.3±3.3 70.8±0.3
sw 56.1±0.5 42.4±0.2 42.9±0.1 70.7±3.0 71.7±0.3 69.9±0.9 42.5±0.1 43.1±0.2 67.2±3.9 71.1±0.6
ar 55.9±0.1 44.2±0.2 44.4±0.0 67.7±1.9 76.9±0.2 67.2±0.2 44.2±0.0 44.3±0.2 63.9±3.8 76.2±0.4
fr 59.2±0.3 48.9±0.1 49.2±0.4 70.5±1.9 75.9±0.2 68.7±0.1 48.9±0.1 49.1±0.5 68.1±2.9 75.9±0.1
ru 58.2±2.0 45.9±0.1 46.2±0.2 70.5±1.9 75.0±0.6 68.0±0.8 45.8±0.0 46.7±0.3 67.2±3.4 75.1±0.1
@@
$\mathbf{r}$ 2 4
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 49.3±0.1 49.3±0.1 49.1±0.1 54.4±0.4 54.0±0.6 49.9±0.1 49.3±0.1 49.3±0.2 54.7±0.4 55.3±0.3
hi 45.9±0.2 45.3±0.1 45.1±0.2 49.9±0.2 51.1±0.3 47.1±0.1 45.2±0.2 45.1±0.1 50.9±0.7 50.6±0.9
es 51.1±0.1 50.3±0.2 50.2±0.0 56.8±0.3 57.7±0.1 52.3±0.3 50.1±0.2 50.0±0.2 57.3±0.3 58.2±0.2
el 48.5±0.2 47.8±0.2 48.0±0.2 51.4±0.3 51.1±0.7 49.7±0.2 47.9±0.2 48.2±0.2 51.5±0.3 51.3±0.4
vi 48.5±0.2 47.3±0.1 47.1±0.2 54.0±0.7 54.5±0.2 49.9±0.2 47.3±0.1 47.1±0.1 54.7±1.1 54.6±0.3
tr 45.3±0.2 44.6±0.0 44.7±0.1 48.3±0.5 49.7±0.1 46.5±0.3 44.6±0.0 44.7±0.1 47.7±0.2 48.8±0.3
de 49.8±0.2 49.3±0.3 49.1±0.2 54.0±0.6 53.3±0.7 50.3±0.1 49.3±0.2 49.4±0.0 54.2±0.4 52.7±0.6
ur 46.9±0.3 46.6±0.2 46.5±0.2 50.1±0.4 53.1±1.0 48.2±0.2 46.5±0.1 46.4±0.2 50.9±0.2 53.5±0.5
en 55.4±0.3 54.9±0.2 54.7±0.1 59.9±0.4 59.6±0.3 56.5±0.6 54.7±0.2 54.9±0.1 59.8±0.6 59.1±0.7
zh 49.3±0.2 48.7±0.2 48.7±0.2 52.5±0.2 51.5±0.4 49.8±0.2 48.6±0.2 48.9±0.1 52.7±0.1 51.5±0.3
th 44.4±0.0 44.7±0.1 44.8±0.2 48.0±0.3 48.9±0.2 45.0±0.4 44.6±0.2 44.6±0.0 47.9±0.5 48.5±0.4
sw 41.2±0.2 40.1±0.1 40.1±0.1 44.5±0.3 46.5±0.7 41.4±0.0 40.1±0.2 40.1±0.1 44.8±0.0 47.2±0.6
ar 46.9±0.1 46.7±0.1 46.6±0.2 50.3±0.4 49.0±0.4 47.9±0.1 46.5±0.1 46.6±0.0 50.2±0.3 49.0±0.2
fr 48.3±0.1 48.7±0.1 48.7±0.1 53.0±0.3 56.3±0.7 48.7±0.1 48.8±0.0 48.5±0.1 53.5±0.5 56.7±0.7
ru 45.1±0.1 44.5±0.1 44.3±0.1 49.1±0.7 52.5±0.4 45.8±0.2 44.5±0.3 44.4±0.0 49.5±0.9 52.1±0.4
$\mathbf{r}$ 8 16
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 51.0±0.3 49.2±0.2 49.3±0.1 54.1±0.2 53.4±0.0 52.7±0.2 49.1±0.1 49.3±0.2 52.5±1.2 53.7±0.5
hi 48.1±0.5 45.4±0.2 45.1±0.2 49.6±1.1 50.6±0.3 48.8±0.0 45.2±0.0 45.3±0.2 49.0±0.3 49.9±0.5
es 54.0±0.3 50.1±0.2 50.3±0.1 57.5±0.8 57.7±0.6 55.5±0.2 50.2±0.2 50.2±0.0 56.4±1.4 57.1±0.5
el 50.9±0.3 48.1±0.1 47.9±0.1 50.9±0.5 50.5±0.3 51.1±1.1 47.9±0.2 47.9±0.1 50.5±1.0 49.9±0.2
vi 51.2±0.3 47.1±0.1 47.2±0.2 53.2±0.3 55.1±0.4 53.5±0.4 46.9±0.2 47.5±0.2 52.5±0.7 54.7±0.9
tr 46.9±0.2 44.7±0.1 44.7±0.1 47.6±0.4 48.0±0.8 48.9±0.8 44.5±0.1 44.8±0.2 47.8±0.3 47.7±0.4
de 52.1±1.1 49.1±0.1 49.3±0.2 54.2±0.2 52.5±0.3 54.4±0.9 49.3±0.2 49.4±0.3 53.9±0.4 52.3±0.4
ur 48.3±0.2 46.3±0.1 46.3±0.2 49.9±0.1 51.9±0.2 49.4±1.0 46.5±0.1 46.5±0.1 48.8±0.8 52.3±0.5
en 58.3±0.2 54.9±0.2 54.9±0.1 59.5±0.6 57.0±1.0 58.9±1.1 54.8±0.2 55.3±0.1 59.1±0.4 57.9±0.7
zh 50.1±0.2 48.5±0.3 48.5±0.1 52.3±0.9 50.7±0.1 51.8±0.5 48.7±0.1 48.9±0.3 51.7±0.8 50.3±0.7
th 45.8±0.2 44.6±0.0 44.5±0.1 47.5±0.2 47.6±0.0 46.9±0.1 44.7±0.2 44.5±0.1 47.1±0.8 47.6±0.3
sw 42.4±0.0 40.1±0.1 40.3±0.1 43.6±0.7 46.7±0.5 43.6±1.0 40.1±0.1 40.2±0.3 43.0±1.1 46.2±0.2
ar 48.1±1.1 46.7±0.1 46.7±0.1 50.3±0.5 48.5±0.7 49.1±1.1 46.7±0.2 46.5±0.1 49.3±0.8 49.0±0.5
fr 50.9±0.1 48.7±0.2 48.5±0.1 52.6±0.6 55.0±0.3 52.1±0.8 48.7±0.1 48.7±0.2 52.1±1.2 54.6±0.7
ru 46.8±0.6 44.3±0.2 44.5±0.4 48.8±0.6 50.7±0.8 48.9±2.0 44.4±0.3 44.7±0.2 48.3±0.7 49.7±0.6